OpenStudy (anonymous):

Simplify the expression. 22+(3^2 –4^2)

OpenStudy (anonymous):

3^2 = ?

OpenStudy (anonymous):

3 TO SECOND POWER

OpenStudy (anonymous):

I DONT KNOW HOW TO MAKE THE SIGNS

OpenStudy (anonymous):

how much it is lol

OpenStudy (anonymous):

6

OpenStudy (anonymous):

nah, 3^2 = 3 into 3 = 9

OpenStudy (anonymous):

I DONT KNOW HOW TO DO THESE AT ALL

OpenStudy (anonymous):

OHH YOU MULTIPLY

OpenStudy (anonymous):

4^2 = ?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

(3^2 –4^2) = 9 - 16 = ?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

cool, 22+(3^2 –4^2) = 22 - 7 = ?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

SO 22+ -7=15

OpenStudy (anonymous):

brilliant, u know how to do it

OpenStudy (anonymous):

THANKS I JUST DONT KNOW THE STEPS I THINK I HAVE TO BE REFRESHED

OpenStudy (anonymous):

u r doin good at it, you are welcome

OpenStudy (anonymous):

THANKS I JUST POSTED ANOTHER IF U FEEL LIKE HELPING ME